OpenStudy (samara89):

A ball has a diameter of 9 in. It consists of 2 parts. The inside is a spherical core with a diameter of 6 in. Surrounding the core is a layer of polyurethane. What is the volume of the polyurethane? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (samara89):

@Igreen

OpenStudy (samara89):

@iGreen

OpenStudy (welshfella):

Volume of polyurethane = volume of the ball - volume of the inner core

OpenStudy (samara89):

@josephinevessey

OpenStudy (samara89):

I do not understand how you are suppose to solve it

OpenStudy (welshfella):

these are both spheres whose volume is (4/3) pi r^3 where r = radius

OpenStudy (samara89):

for 9 we have to divide by 2 right since it is diameter

OpenStudy (samara89):

and change it into radius

OpenStudy (welshfella):

use pi = 3.14 as in the question and the radius of each sphere = ha;lf the diameter

OpenStudy (josephinevessey):

i dont know how to solve it srry

OpenStudy (samara89):

9 divided by 2

OpenStudy (welshfella):

yes the radius of the whole ball = 4.5 and radius of inner core = 1/2 * 6 = 3

OpenStudy (samara89):

what do we do now?

OpenStudy (welshfella):

plug those values into the formulas for the volume and work it out volume = (4/3)*3.14*4.5^3 - (4/3)*3.14*3^3

OpenStudy (welshfella):

use your calculator

OpenStudy (samara89):

first one I got 381.51

OpenStudy (welshfella):

yes thats right - that is the volume of the whole ball

OpenStudy (samara89):

second one I got 113.04

OpenStudy (samara89):

what now?

OpenStudy (welshfella):

well look at the formula - you subtract them

OpenStudy (samara89):

is the answer 268.47

OpenStudy (welshfella):

= (4/3)*3.14*4.5^3 - (4/3)*3.14*3^3

OpenStudy (welshfella):

yes

OpenStudy (samara89):

I have another question

OpenStudy (welshfella):

plz post it seperately